OpenStudy (anonymous):

Which line shows the first error in the solution? (look in comments for problem) A. Line 2 B. Line 3 C. Line 4 D. Line 5

OpenStudy (anonymous):

(1) 5(x-5)=15 (2) 5x-25=15 (3) 5x-25=15-25 (4) 5x= -10 (5) \[\frac{ 5x }{ 5 } = \frac{ -10 }{ 5 }\] (6) x= -2

OpenStudy (shamil98):

Line 3.

OpenStudy (anonymous):

Line 3. 25 is subtracted from the right side, but not from the left.

OpenStudy (shamil98):

You are supposed to add the 25 to both sides not subtract.

OpenStudy (anonymous):

oh okay. thx :)